OpenStudy (anonymous):

find sum to n terms of the series 1.2+2.2^2+3.2^3+.... to n terms

OpenStudy (anonymous):

\[\sum_{i=1}^{n}i \cdot2^i\]

OpenStudy (anonymous):

the answer given is 2+ (n-1) 2^(n+1) ... how to get this ??

OpenStudy (anonymous):

is it 1.2 2.2^2 3.2^3 or 1, 2^2, 3^3 ???

OpenStudy (anonymous):

it is 1.2 2.2^2 3.2^3

OpenStudy (anonymous):

\[\sum_{i=1}^{n}i \cdot 2^i=2^n \sum_{i=1}^{n}\frac{i}{2^{n-i}}\] and ehuman, it's not an arithmetic series

OpenStudy (anonymous):

geometric my bad

OpenStudy (anonymous):

not quite geometric either

OpenStudy (anonymous):

Sum of series is \[Sn=n/2*(2a(n-1)d)\]

OpenStudy (anonymous):

it is a arithmetico - geometric series .... yes i know the rules, somehow not getting the answer.. the answer given is 2+ (n-1) 2^(n+1) ...so if you post the solution completely i will understand how you did it

OpenStudy (anonymous):

close though

OpenStudy (anonymous):

Let \(\large S_n = 1\cdot 2 + 2\cdot 2^2 + 3\cdot 2^3 +\ldots +n\cdot 2^n\). Then it follows that \(\large 2S_n = 1\cdot 2^2+2\cdot 2^3+3\cdot 2^4+\ldots +n\cdot 2^{n+1}\) Now consider \(\large 2S_n - S_n\) and simplify that difference. You should get the formula you're looking for. Can you take things from here? :-)

OpenStudy (anonymous):

@ChristopherToni : thanks a lot .. i got the answer :)